Jacob has a certain amount of fencing he wants to use to make a dog run. If he makes the length 36 feet, the width will be 9 feet, but if he makes the length 40 feet, the width will be only 5 feet. Write an equation in the form y = mx + b that can be used to find the width, y, for any given length x.
Answer by stanbon(75887)

Jacob has a certain amount of fencing he wants to use to make a dog run. If he makes the length 36 feet, the width will be 9 feet, but if he makes the length 40 feet, the width will be only 5 feet. Write an equation in the form y = mx + b that can be used to find the width, y, for any given length x.
-----
You have two points relating width to length::
(9,36) and (5,40)
-----
slope = (40-36)/(5-9) = 4/-4 = -1
-----
Form:: y = mx + b
Solve for "b":
40 = -1*5 + b
b = 45
-----
Equation:
y = -x + 45
------------
Cheers,
Stan H.
------------